A train traveling at an average speed of 100 km /h covers a certain distance in 3 hours 36. At what avarage speed must the train travel to cover the same distance in 2 hours 30 minutes?
Answer by ewatrrr(24785)


Hi
 D = rt 
D (100kmi/h)(216/60)h = 360mi

r = D/t = 360mi/(2.5h) = 144mi/h  
144 mi/h is the average speed to cover the same distance in 2.5hr. 

Wish You the Best in your Studies.